# A Survey on 3D Gaussian Splatting

### v4

@@ Abstract
3D Gaussian splatting (GS) has recently emerged as a transformative technique in the realm of explicit radiance field and computer graphics. This innovative approach, characterized by the utilization of millions of learnable 3D Gaussians, represents a significant departure from mainstream neural radiance field approaches, which predominantly use implicit, coordinate-based models to map spatial coordinates to pixel values. 3D GS, with its explicit scene representation and differentiable rendering algorithm, not only promises real-time rendering capability but also introduces unprecedented levels of editability. This positions 3D GS as a potential game-changer for the next generation of 3D reconstruction and representation. In the present paper, we provide the first systematic overview of the recent developments and critical contributions in the domain of 3D GS. We begin with a detailed exploration of the underlying principles and the driving forces behind the emergence of 3D GS, laying the groundwork for understanding its significance. A focal point of our discussion is the practical applicability of 3D GS. By enabling unprecedented rendering speed, 3D GS opens up a plethora of applications, ranging from virtual reality to interactive media and beyond. This is complemented by a comparative analysis of leading 3D GS models, evaluated across various benchmark tasks to highlight their performance and practical utility. The survey concludes by identifying current challenges and suggesting potential avenues for future research in this domain. Through this survey, we aim to provide a valuable resource for both newcomers and seasoned researchers, fostering further exploration and advancement in applicable and explicit radiance field representation.
@@ 1 Introduction
The objective of image based 3D scene reconstruction is to convert a collection of views or videos capturing a scene into a 3D model that can be processed and understood by computers. This hard and long-standing problem is fundamental for machines to comprehend the complexity of real-world environments, facilitating a wide array of applications such as 3D modeling and animation, robot navigation, historical preservation, augmented/virtual reality, and autonomous driving.
The journey of 3D scene reconstruction began long before the surge of deep learning, with early endeavors focusing on light fields and basic scene reconstruction methods [1, 2, 3]. These early attempts, however, were limited by their reliance on dense sampling and structured capture, leading to significant challenges in handling complex scenes and lighting conditions. The emergence of structure-from-motion [4] and subsequent advancements in multi-view stereo [5] algorithms provided a more robust framework for 3D scene reconstruction. Despite these advancements, such methods struggled with novel-view synthesis and lacked compatibility with deep scene understanding models. NeRF represents a quantum leap in this progression. By leveraging deep neural networks, NeRF enabled the direct mapping of spatial coordinates to color and density. The success of NeRF hinged on its ability to create continuous, volumetric scene functions, producing results with unprecedented detail and realism. However, as with any burgeoning technology, this implementation came at a cost: i) Computational Intensity. NeRF based methods are computationally intensive [6, 7, 8, 9], often requiring extensive training times and substantial resources for rendering, especially for high-resolution outputs. ii) Editability. Manipulating scenes represented implicitly is challenging, since direct modifications to the neural network’s weights are not intuitively related to changes in geometric or appearance properties of the scene.
Figure 1: The number of follow-up arXiv papers and official github stars on 3D GS is increasing every month.
@@
In an effort to assist readers in keeping pace with the swift evolution of 3D GS, we provide the first survey on 3D GS, which presents a systematic and timely collection of the most significant literature on the topic. Given that 3D GS is a very recent innovation (Fig. 1), this survey focuses in particular on its principles, and the diverse developments and contributions that have emerged since its introduction. The selected follow-up works are primarily sourced from top-tier conferences, to provide a thorough and up-to-date analysis of the theoretical foundations, remarkable developments, and burgeoning applications of 3D GS. Acknowledging the nascent yet rapidly evolving nature of 3D GS, this survey is inevitably a biased view, but we strive to offer a balanced perspective that reflects both the current state and the future potential of this field. Our aim is to encapsulate the primary research trends and serve as a valuable resource for both researchers and practitioners eager to understand and contribute to this rapidly evolving domain.
@@ 2 Background
In this section, we first provide a brief formulation of radiance fields (Sec. 2.1), a key concept in scene rendering. It outlines two main types of radiance field representations: implicit, like NeRF [12], which uses neural networks for a straightforward yet computationally demanding rendering; and explicit, like grid [25], which employs discrete structures for quicker access but at the cost of higher memory use. Sec. 2.2 further establishes linkages with relevant fields such as scene reconstruction and rendering. For a comprehensive overview of radiance fields, scene reconstruction and representation, and rendering methods, please see the excellent surveys [26, 27, 28, 29, 30] for more insights.
2.1 Problem Formulation
2.1.1 Radiance Field
A radiance field is a representation of light distribution in a 3D space, which captures how light interacts with surfaces and materials in the environment [27]. Mathematically, a radiance field can be described as a function $L\!:\!\mathbb{R}^{5}\!\mapsto\!\mathbb{R}^{+}$, where $L(x,y,z,\theta,\phi)$ maps a point in space $(x,y,z)$, and a direction specified by spherical coordinates $(\theta,\phi)$, to a non-negative radiance value. Radiance fields can be encapsulated through implicit or explicit representations, each with specific advantages for scene representation and rendering.
2.1.2 Implicit Radiance Field
@@
An implicit radiance field represents light distribution in a scene without explicitly defining the geometry of the scene. In the deep learning era, it often uses neural networks to learn a continuous volumetric scene representation [31, 32]. The most prominent example is NeRF [12]. In NeRF, a neural network, typically a multi-layer perceptron (MLP), is used to map a set of spatial coordinates $(x,y,z)$ and viewing directions $(\theta,\phi)$ to color and density values. The radiance at any point is not stored explicitly but is computed on-the-fly by querying the MLP. Hence, the function can be written as:
$L_{\text{implicit}}(x,y,z,\theta,\phi)=\text{MLP}(x,y,z,\theta,\phi).$ (1)
@@
This format allows for a differentiable and compact representation of complex scenes, albeit often at the cost of high computational load due to volumetric ray marching [10].
2.1.3 Explicit Radiance Field
@@
In contrast, an explicit radiance field directly represents the distribution of light in a discrete spatial structure, such as a voxel grid or a set of points [25, 33]. Each element in this structure stores the radiance information for its respective location in space. This allows for more direct and often faster access to radiance data but at the cost of higher memory usage and potentially lower resolution. A generic form for an explicit radiance field representation can be written as:
$L_{\text{explicit}}(x,y,z,\theta,\phi)=\text{DataStructure}[(x,y,z)]\cdot f(\theta,\phi),$ (2)
where DataStructure could be in the format of volumes, point clouds, etc. $f(\theta,\phi)$ is a function that modifies the radiance based on the viewing direction.
2.1.4 3D Gaussian Splatting: Best-of-Both Worlds
@@
3D GS [10] is an explicit radiance field with the advantages of implicit radiance fields. Concretely, it leverages the strengths of both paradigms by utilizing learnable 3D Gaussians as a flexible and efficient representation. These Gaussians are optimized under the supervision of multi-view images to accurately represent the scene. Such a 3D Gaussian based differentiable pipeline combines the benefits of neural network-based optimization and explicit, structured data storage. This hybrid approach aims to achieve real-time, high-quality rendering and requires less training time, particularly for complex scenes and high-resolution outputs. The 3D Gaussian representation is formulated as:
$L_{\text{3DGS}}(x,y,z,\theta,\phi)=\sum_{i}G(x,y,z,\bm{\mu}_{i},\bm{\Sigma}_{i})\cdot c_{i}(\theta,\phi),$ (3)
where $G$ is the Gaussian function with mean $\bm{\mu}_{i}$ and covariance $\bm{\Sigma}_{i}$, and $c$ represents the view-dependent color.
@@ 2.2 Context and Terminology
Several techniques and research disciplines possess a close relationship with 3D GS, which will be described briefly in the following sections for clarity.
2.2.1 Scene Reconstruction and Rendering
Roughly speaking, scene reconstruction involves creating a 3D model of a scene from a collection of images or other data. Rendering is a more specific term that focuses on transforming computer-readable information (e.g., 3D objects in the scene) to pixel-based images. Early techniques generated realistic images based on the light fields [1, 2, 3]. The structure-from-motion [4] and multi-view stereo [5] algorithms further advanced this field by estimating 3D structures from image sequences. These historical methods provide a solid foundation for more complex scene reconstruction and rendering techniques [34, 35, 36, 37].
2.2.2 Neural Rendering and Radiance Fields
Neural rendering integrates deep learning with traditional graphics techniques to create photorealistic images. Early attempts utilized convolutional networks for estimating blending weights [36] or texture-space solutions [38]. As mentioned in Sec. 2.1.1, the radiance field represents a function that describes the amount of light traveling in every direction through every point in space. NeRFs [12, 9, 8] use neural networks, typically MLPs, to model the radiance fields, enabling detailed and realistic scene rendering.
2.2.3 Volumetric Representations and Ray-Marching
Volumetric representations model objects and scenes not just as surfaces but as volumes filled with materials or empty space. This allows for a more accurate rendering of phenomena like fog, smoke, or translucent materials. Ray-marching is a technique used with volumetric representations to render images by tracing the path of light through a volume [11]. NeRF [12] shares the same spirit of volumetric ray-marching and introduces importance sampling and positional encoding to improve the quality of synthesized images. While providing high-quality results, volumetric ray-marching is computationally expensive, motivating the search for more efficient methods like 3D GS.
2.2.4 Point-based Rendering
Point-based rendering is a technique for visualizing 3D scenes using points rather than traditional polygons. This method is particularly effective for rendering complex, unstructured, or sparse geometric data. Points can be augmented with additional properties like learnable neural descriptors [39, 40], and rendered efficiently [41, 42], but this approach suffers from issues like holes in the rendering or aliasing effects. 3D GS [10] extends this concept by using anisotropic Gaussians for a more continuous and cohesive representation of the scene. More implementation details will be further discussed in Sec. 3.
@@ 2.2.4 Point-based Rendering
$x,y,z,\theta,\phi$
$c~{}\&~{}\alpha$
Figure 3: NeRFs vs. 3D GS. (a) NeRF samples along the ray and then queries the MLP to obtain corresponding colors and opacities, which can be seen as a backward mapping (ray tracing). (b) In contrast, 3D GS projects all 3D Gaussians into the image space (i.e., splatting) and then performs parallel rendering, which can be viewed as a forward mapping (splatting and rasterization). Best viewed in color.
@@ 3.1 Rendering with Learned 3D Gaussians
Consider a scene represented by (millions of) optimized 3D Gaussians. The objective is to generate an image from a specified camera pose. Recall that NeRFs approach this task through computationally demanding volumetric ray-marching, sampling 3D space points per pixel. Such a paradigm struggles with high-resolution image synthesis, failing to achieve real-time rendering, especially for platforms with limited computing resources [10]. By contrast, 3D GS begins by projecting these 3D Gaussians onto a pixel-based image plane, a process termed “splatting” (see Fig. 3b). Afterwards, 3D GS sorts these Gaussians and computes the value for each pixel. As shown in Fig. 3, the rendering of NeRFs and 3D GS can be viewed as an inverse process of each other. In what follows, we begin with the definition of a 3D Gaussian, which is the minimal element of the scene representation in 3D GS. Next, we describe how these 3D Gaussians can be used for differentiable rendering. Finally, we introduce the acceleration technique used in 3D GS, which is the key to fast rendering.
@@
$\bullet$ Frustum Culling. Given a specified camera pose, this step determines which 3D Gaussians are outside the camera’s frustum. By doing so, 3D Gaussians outside the given view will not be involved in the subsequent computation, thus saving computational resources.
$\bullet$ Splatting. In this step, 3D Gaussians (ellipsoids) in 3D space are projected into the 2D image space (ellipses) for rendering. Given the viewing transformation $\bm{W}$ and 3D covariance matrix $\bm{\Sigma}$, the projected 2D covariance matrix $\bm{\Sigma}^{\prime}$ is computed using:
@@
where $\bm{J}$ is the Jacobian of the affine approximation of the projective transformation [43, 10].
@@
$\bullet$ Rendering by Pixels. Before delving into the final version of 3D GS which utilizes several techniques to boost parallel computation, we first elaborate on its simpler form to offer insights into its basic working mechanism. Given the position of a pixel $\bm{x}$, its distance to all overlapping Gaussians, i.e., the depths of these Gaussians, can be computed through the viewing transformation $\bm{W}$, forming a sorted list of Gaussians $\mathcal{N}$. Then, alpha compositing is adopted to compute the final color of this pixel:
@@
5
$\bullet$ Tiles (Patches). To avoid the cost computation of deriving Gaussians for each pixel, 3D GS shifts the precision from pixel-level to patch-level detail. Concretely, 3D GS initially divides the image into multiple non-overlapping patches (called “tiles” in the original paper [10]). Fig. 4b provides an illustration of tiles. Each tile comprises 16$\times$16 pixels as suggested in [10]. 3D GS further determines which tiles intersect with these projected Gaussians. Given that a projected Gaussian may cover several tiles, a logical method involves replicating the Gaussian, assigning each copy an identifier (i.e., a tile ID) for the relevant tile.

### v5

@@ Abstract
3D Gaussian splatting (GS) has emerged as a transformative technique in explicit radiance field and computer graphics. This innovative approach, characterized by the use of millions of learnable 3D Gaussians, represents a significant departure from mainstream neural radiance field approaches, which predominantly use implicit, coordinate-based models to map spatial coordinates to pixel values. 3D GS, with its explicit scene representation and differentiable rendering algorithm, not only promises real-time rendering capability but also introduces unprecedented levels of editability. This positions 3D GS as a potential game-changer for the next generation of 3D reconstruction and representation. In the present paper, we provide the first systematic overview of the recent developments and critical contributions in the domain of 3D GS. We begin with a detailed exploration of the underlying principles and the driving forces behind the emergence of 3D GS, laying the groundwork for understanding its significance. A focal point of our discussion is the practical applicability of 3D GS. By enabling unprecedented rendering speed, 3D GS opens up a plethora of applications, ranging from virtual reality to interactive media and beyond. This is complemented by a comparative analysis of leading 3D GS models, evaluated across various benchmark tasks to highlight their performance and practical utility. The survey concludes by identifying current challenges and suggesting potential avenues for future research. Through this survey, we aim to provide a valuable resource for both newcomers and seasoned researchers, fostering further exploration and advancement in explicit radiance field.
@@ 1 Introduction
The objective of image based 3D scene reconstruction is to convert a collection of views or videos capturing a scene into a digital 3D model that can be computationally processed, analyzed, and manipulated. This hard and long-standing problem is fundamental for machines to comprehend the complexity of real-world environments, facilitating a wide array of applications such as 3D modeling and animation, robot navigation, historical preservation, augmented/virtual reality, and autonomous driving.
The journey of 3D scene reconstruction began long before the surge of deep learning, with early endeavors focusing on light fields and basic scene reconstruction methods [1, 2, 3]. These early attempts, however, were limited by their reliance on dense sampling and structured capture, leading to significant challenges in handling complex scenes and lighting conditions. The emergence of structure-from-motion [4] and subsequent advancements in multi-view stereo [5] algorithms provided a more robust framework for 3D scene reconstruction. Despite these advancements, such methods struggled with novel-view synthesis and texture loss. NeRF represents a quantum leap in this progression. By leveraging deep neural networks, NeRF enabled the direct mapping of spatial coordinates to color and density. The success of NeRF hinged on its ability to create continuous, volumetric scene functions, producing results with unprecedented fidelity. However, as with any burgeoning technology, this implementation came at a cost: i) Computational Intensity. NeRF based methods are computationally intensive [6, 7, 8, 9], often requiring extensive training times and substantial resources for rendering, especially for high-resolution outputs. ii) Editability. Manipulating scenes represented implicitly is challenging, since direct modifications to the neural network’s weights are not intuitively related to changes in geometric or appearance properties of the scene.
Figure 1: The number of published papers and official GitHub stars on 3D GS. The set of statistics is sourced from # Papers and # GitHub Stars.
@@
In an effort to assist readers in keeping pace with the swift evolution of 3D GS, we provide the first survey on 3D GS, which presents a systematic and timely collection of the most significant literature on the topic. Given that 3D GS is a very recent innovation (Fig. 1), this survey focuses in particular on its principles, and the diverse developments and contributions that have emerged since its introduction. The selected follow-up works are primarily sourced from top-tier conferences, to provide a thorough and up-to-date (Dec. 2024) analysis of the theoretical foundations, remarkable developments, and burgeoning applications of 3D GS. Acknowledging the nascent yet rapidly evolving nature of 3D GS, this survey is inevitably a biased view, but we strive to offer a balanced perspective that reflects both the current state and the future potential of this field. Our aim is to encapsulate the primary research trends and serve as a valuable resource for both researchers and practitioners eager to understand and contribute to this rapidly evolving domain. The distinctions of this survey from existing literature [25, 26, 27, 28] are evident in the following aspects:
$\bullet$ We provide the first systematic and comprehensive review that examines 3D GS from a macro-level perspective by establishing clear taxonomies and frameworks. This high-level systematization helps researchers identify trends and potential directions that might not be apparent from paper-specific reviews. Our organizational structure serves as a roadmap for understanding how different approaches relate to and build upon each other within the 3D GS ecosystem.
$\bullet$ This paper is the first and only survey to thoroughly delve into the theoretical background and fundamental principles of 3D GS. The comprehensive coverage makes the field more approachable for newcomers while providing valuable insights for experienced researchers.
$\bullet$ To ensure our survey remains relevant and offer long-term value in this rapidly evolving field, we maintain two dynamic GitHub repositories: one that follows our survey’s organizational structure and another that includes comprehensive performance comparisons with analysis data.
@@ 2 Background
In this section, we first provide a brief formulation of radiance fields (Sec. 2.1), including both implicit and explicit ones. Sec. 2.2 further establishes linkages with relevant rendering algorithms and terminologies. For a comprehensive overview of radiance fields, scene reconstruction and representation, and rendering methods, please see the excellent surveys [29, 30, 31, 32, 33] for more insights.
2.1 Radiance Field
@@
$\bullet$ Implicit Radiance Field. An implicit radiance field represents light distribution in a scene without explicitly defining the geometry of the scene. In the deep learning era, neural networks are often used to learn a continuous volumetric scene representation [34, 35]. The most prominent example is NeRF [12]. In NeRF (Fig. 3a), one or more MLPs are used to map a set of spatial coordinates $(x,y,z)$ and viewing directions $(\theta,\phi)$ to color $c$ and volume density $\sigma$:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(c,\sigma)\leftarrow\text{MLP}(x,y,z,\theta,\phi)}.$ (1)
@@
This format allows for a differentiable and compact representation of complex scenes, albeit often at the cost of high computational load due to volumetric ray marching. Note that typically, the color $c$ is direction-dependent, whereas the volume density $\sigma$ is not [12].
@@
$\bullet$ Explicit Radiance Field. An explicit radiance field directly represents the distribution of light in a discrete spatial structure, such as a voxel grid or a set of points [36, 37]. Each element in this structure stores the radiance information for its respective location. This allows for direct and often faster access to radiance data but at the cost of higher memory usage and potentially lower resolution. Similar to the implicit radiance field, the explicit one is written as:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(c,\sigma)\leftarrow\text{DataStructure}(x,y,z,\theta,\phi)},$ (2)
where DataStructure could be in the format of volumes, point clouds, etc. DataStructure encodes directional color in two main ways. One is encoding high-dimensional features that are subsequently decoded by a lightweight MLP. Another one is directly storing coefficients of directional basis functions, such as spherical harmonics or spherical Gaussians, where the final color is computed as a function of these coefficients and the viewing direction.
@@
$\bullet$ 3D Gaussian Splatting: Best-of-Both Worlds. 3D GS [10] is an explicit radiance field with the advantages of implicit radiance fields. Concretely, it leverages the strengths of both paradigms by utilizing learnable 3D Gaussians as the basis elements of DataStructure. Note that 3D GS encodes the opacity $\alpha$ directly for each Gaussian, as opposed to approaches of first establishing density $\sigma$ and then computing opacity based on that density. As in previous reconstruction work, 3D Gaussians are optimized under the supervision of multi-view images to represent the scene. Such a 3D Gaussian based differentiable pipeline combines the benefits of neural network based optimization and explicit, structured data storage. This hybrid approach aims to achieve real-time, high-quality rendering and requires less training time, particularly for complex scenes and high-resolution outputs.
@@ 2.2 Context and Terminology
$\bullet$ Volumetric rendering aims to transform a 3D volumetric representation into an image by integrating radiance along camera rays. A camera ray $\bm{r}(t)$ can be parameterized as: $\bm{r}(t)\!=\!\bm{o}\!+\!t\bm{d},t\!\in\![t_{\text{near}},t_{\text{far}}],$ where $\bm{o}$ represents the ray origin (camera center), $\bm{d}$ is the ray direction, and $t$ indicates the distance along the ray between near and far clipping planes. The pixel color $C(\bm{r})$ is computed through a line integral along the ray $\bm{r}(t)$, mathematically expressed as [12]:
[…]
where $\sigma(\bm{r}(t))$ is the volume density at point $\bm{r}(t)$, $c(\bm{r}(t),\bm{d})$ is the color at that point, and $T(t)$ is the transmittance. Ray-marching directly approximates the volumetric rendering integral by systematically “stepping” along a ray and sampling the scene’s properties at discrete intervals. NeRF [12] shares the same spirit of ray-marching and introduces importance sampling and positional encoding to improve the quality of synthesized images. While providing high-quality results, ray-marching is computationally expensive, especially for high-resolution images.
$\bullet$ Point-based rendering represents another class of rendering algorithms, of which 3D GS introduces a notable implementation. Its simplest form [38] rasterizes point clouds with a fixed size, which introduces drawbacks such as holes and rendering artifacts. Seminal works addressed these limitations through various methods, including: i) splatting point primitives with a spatial extent [14, 39, 40, 15], and ii) more recently, embedding neural features directly into points for subsequent network-based rendering [41, 42]. 3D GS uses 3D Gaussian as the point primitive that contains explicit attributes (e.g., color and opacity) instead of implicit neural features. The rendering approach, i.e., point-based $\alpha$-blending (exemplified in Eq. 5), shares the same image formation model as NeRF-style volumetric rendering (Eq. 3[10], but demonstrates substantial speed advantages. This advantage originates from fundamental algorithmic differences. NeRFs approximate a line integral along a ray for each pixel, requiring expensive sampling. Point-based methods render point clouds using rasterization, which inherently benefits from parallel computational strategies [43].
@@ 3.1 Rendering with Learned 3D Gaussians
Consider a scene represented by (millions of) optimized 3D Gaussians. The objective is to generate an image from a specified camera pose. Recall that NeRFs approach this task through computationally demanding volumetric ray-marching, sampling 3D space points per pixel. Such a paradigm struggles with high-resolution image synthesis, failing to achieve real-time rendering, especially for platforms with limited computing resources [10]. By contrast, 3D GS begins by projecting these 3D Gaussians onto a pixel-based image plane, a process termed “splatting” [39, 40] (see Fig. 3b). Afterwards, 3D GS sorts these Gaussians and computes the value for each pixel. As shown in Fig. 3, the rendering of NeRFs and 3D GS can be viewed as an inverse process of each other. In what follows, we begin with the definition of a 3D Gaussian, which is the minimal element of the scene representation in 3D GS. Next, we describe how these 3D Gaussians can be used for differentiable rendering. Finally, we introduce the acceleration technique used in 3D GS, which is the key to fast rendering.
@@
$\bullet$ Frustum Culling. Given a specified camera pose, this step determines which 3D Gaussians are outside the camera’s frustum. By doing so, 3D Gaussians outside the given view will not be involved in the subsequent computation.
$\bullet$ Splatting. In this step, 3D Gaussians (ellipsoids) in 3D space are projected into 2D image space (ellipses). The projection proceeds through two transformations: first, transforming 3D Gaussians from world coordinates to camera coordinates using the viewing transformation, and subsequently splatting these Gaussians into 2D image space via an approximation of the projective transformation. Mathematically, given the 3D covariance matrix $\bm{\Sigma}$ describing a 3D Gaussian’s spatial distribution, and the viewing transformation matrix $\bm{W}$, the 2D covariance matrix $\bm{\Sigma}^{\prime}$ characterizing the projected 2D Gaussian is computed through:
@@
where $\bm{J}$ is the Jacobian of the affine approximation of the projective transformation [39, 10]. One might wonder why the standard camera intrinsics based projective transformation is not used here. This is because its mappings are not affine and therefore cannot directly project $\bm{\Sigma}$. 3D GS adopts an affine one proposed in [39] which approximates the projective transformation using the first two terms (including $\bm{J}$) of the Taylor expansion (see Sec. 4.4 in [39]).
$x,y,z,\theta,\phi$
$c~{}\&~{}\sigma$
Figure 3: NeRFs vs. 3D GS. (a) NeRF samples along the ray and then queries the MLP to obtain corresponding colors and densities, which can be seen as a backward mapping (ray tracing). (b) In contrast, 3D GS projects all 3D Gaussians into the image space (i.e., splatting) and then performs parallel rendering, which can be viewed as a forward mapping (rasterization). Best viewed in color.
@@
$\bullet$ Rendering by Pixels. Before delving into the final version of 3D GS which utilizes several techniques to boost parallel computation, we first elaborate on its simpler form to offer insights into its basic working mechanism. Given the position of a pixel $\bm{x}$, its distance to all overlapping Gaussians, i.e., the depths of these Gaussians, can be computed through the viewing transformation matrix $\bm{W}$, forming a sorted list of Gaussians $\mathcal{N}$. Then, $\alpha$-blending is adopted to compute the final color of this pixel:
@@
$\bullet$ Tiles (Patches). To avoid the cost computation of deriving Gaussians for each pixel, 3D GS shifts the precision from pixel-level to patch-level detail, which is inspired by tile-based rasterization [43]. Concretely, 3D GS initially divides the image into multiple non-overlapping patches (tiles). Fig. 4b provides an illustration of tiles. Each tile comprises 16$\times$16 pixels as suggested in [10]. 3D GS further determines which tiles intersect with these projected Gaussians. Given that a projected Gaussian may cover several tiles, a logical method involves replicating the Gaussian, assigning each copy an identifier (i.e., a tile ID) for the relevant tile.